What is the energy of a photon with a frequency of 7.33 X 10^14 s^-1? (h= 6.626 X 10^-34 J X S)
1
Expert's answer
2020-04-06T12:03:20-0400

E = hf 

h = 6.626*10-34 Js

f = 7.33 * 1014 s-1

E = 6.626*10-34 x 7.33 *1014 = 48.6* 10 -20J